# Multiple Invertible and Partial-Equivariant Function for Latent Vector Transformation to Enhance Disentanglement in VAEs

### v1

@@ Appendix C Proof
If $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{a}$ and $G_{L}^{a}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ by the equivariance of $q_{\phi}$.
@@
Proof. The function $\psi_{G_{S}}(\cdot)$ is an equivariant function over group elements in $G_{S}\cap G_{L}$ by Proposition 3.1. Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$. Let $g_{L}^{a}$ be a group element in $G_{S}\cap G_{L}$, and $g_{I}^{a}$ is a group element in $G_{I}$ corresponding to $G_{S}\cap G_{L}$, and $g_{T}^{a}$ is a group element where corresponding to $G_{S}\cap G_{L}$ on the latent vector space transformed from the original latent vector space. Then, group element $g_{T}^{a}$ is equal to $g_{L}^{a}$:
@@
$\begin{split}&\text{then}~{}g_{L}^{a}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{a}\psi_{G_{S}}({\bm{z}}_{1})~{}(\because\hat{{\bm{z}}}_{2}=g_{T}^{a}\hat{{\bm{z}}}_{1})\\ &\Rightarrow(g_{L}^{a}-g_{T}^{a})\psi_{G_{S}}({\bm{z}}_{1})=\mathbf{0},\end{split}$ (16)
where $\mathbf{0}$ is a zero vector. Eq. 16 is defined when $\forall{\bm{z}}\in\mathcal{Z}$ by the equivariance definition. In other words, Eq. 16 is satisfied only if the kernel (linear algebra) of $g_{L}^{a}-g_{T}^{a}$, notated as ker$(g_{L}^{a}-g_{T}^{a})$, includes the basis of $\mathbb{R}^{n}$ vector space. If the standard basis of $\mathbb{R}^{n}$ vector space is in ker$(g_{L}^{a}-g_{T}^{a})$, then $(g_{L}^{a}-g_{T}^{a})=\mathbf{0}_{n,n}$, where $\mathbf{0}_{n,n}$ is an n by n zero matrix. Other bases of $\mathbb{R}^{n}$ vector space are expressed by the standard basis. Therefore $g_{L}^{a}-g_{T}^{a}=\mathbf{0}_{n,n}$.
@@
Then, $\psi_{G_{S}}(g_{L}^{a}{\bm{z}}_{1})=g_{L}^{a}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{a}\psi_{G_{S}}({\bm{z}}_{1})$. The encoder is an equivariant function over input space $\mathcal{X}$ as $q_{\phi}(g_{I}^{a}{\bm{x}}_{1})=g_{L}^{a}q_{\phi}({\bm{x}}_{1})$. Mixing two equivarience property, we can derive another equivariance relation $g_{T}^{a}\psi_{G_{S}}(q_{\phi}({\bm{x}}_{1}))=\psi_{G_{S}}(q_{\phi}(g_{I}^{a}{\bm{x}}_{1}))$ This result implies that the equivariance between input space and a latent space is preserved for $G_{S}\cap G_{L}$ if the latent vector ${\bm{z}}$ is transformed by $\psi_{G_{S}}.$$\blacksquare$
@@
We show that $\psi_{G_{S}}$ preserves equivariance between $G_{L}^{a}$ and $G_{I}^{a}$. If there exists equivariant function between input and latent vector space, there should be a group $G_{L}$ for a latent space and its corresponding group $G_{I}$ in an input space by definition of equivariance ($q_{\phi}(g_{I}x)=g_{L}q_{\phi}(x)$).
In other words, $\psi_{G_{S}}(\cdot)$ guarantees to preserve the equivariance of I2L-transformation to certain symmetries in $G_{S}\cap G_{L}$ after IPE-transformation as shown in Figure 2.
Let $P(B)$ be the probability of $\psi(\cdot)\in B$ for a subset $B\subset M_{n}(\mathbb{R})$ after VAE training, and $Pr(\psi_{B}\in B^{\prime})$ be the conditional probability of $\psi(\cdot)\in B^{\prime}$ given $\psi(\cdot)\in B$.
Then,
Proposition C.2.
Proof. All $\mathbf{e}^{{\bm{S}}}\in E_{S}$ are in $E_{M}$ since $Sym_{n}(\mathbb{R})\subset M_{n}(\mathbb{R})$. However, $E_{M}\not\subset E_{S}$ because $\mathbf{e}^{{\bm{S}}}$ is always symmetric, but $\mathbf{e}^{{\bm{M}}}$ can be an asymmetric matrix. All elements of $E_{S}$ are symmetric because of the matrix exponential property that $\mathbf{e}^{{\bm{M}}^{\intercal}}=(\mathbf{e}^{{\bm{M}}})^{\intercal}$. If ${\bm{M}}$ is a symmetric matrix then $\mathbf{e}^{{\bm{M}}^{\intercal}}=\mathbf{e}^{{\bm{M}}}=(\mathbf{e}^{{\bm{M}}})^{\intercal}$. Therefore, if ${\bm{M}}$ is symmetric then the exponential of ${\bm{M}}$ is also symmetric. We show a counter example to $E_{M}\subset E_{S}$. When ${\bm{M}}=\begin{bmatrix}1&1\\ 0&1\end{bmatrix}$,
$\begin{split}\mathbf{e}^{{\bm{M}}}&=\sum_{k=0}^{\infty}\frac{1}{k!}{\bm{M}}^{k}\\ &=I+\begin{bmatrix}1&1\\ 0&1\end{bmatrix}+\frac{1}{2!}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}^{2}+\cdots+\frac{1}{(n-1)!}\begin{bmatrix}1&1\\ 0&1\end{bmatrix}^{(n-1)}+\cdots\\ &=I+\begin{bmatrix}\sum_{n=0}^{\infty}\frac{1}{n!}&1+\sum_{n=0}^{\infty}\frac{1}{(n-1)!}\\ 0&\sum_{n=0}^{\infty}\frac{1}{n!}\end{bmatrix}\\ &=\begin{bmatrix}1+e&1+e\\ 0&1+e\end{bmatrix}.\end{split}$ (17)
The matrix $\mathbf{e}^{{\bm{M}}}$ is asymmetric and not in $E_{S}$. Therefore $E_{M}\not\subset E_{S}$. Therefore, the probability $Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{S})}$ is greater than $Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(E_{M})}$. In the same way, $Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})=\frac{P(G_{S})}{P(M_{n}(\mathbb{R}))}$ because $E_{M}\subset M_{n}(\mathbb{R})$ and non-invertible functions are only in $M_{n}(\mathbb{R})$. $\blacksquare$
C.0.1 KL Divergence for Evidence of Lower Bound
@@
The KL divergence of Gaussian distribution Kingma & Welling (2013) is computed using mean and variance, which are the parameters of a Gaussian distribution. To introduce a loss as the KL divergence of Gaussian distribution, we compute KL divergence of the exponential family in Eq. 2 using the learnable parameter $T(\cdot)$ and $A(\cdot)$ with given natural parameter $\bm{\theta}_{\hat{{\bm{z}}}}$ and $\bm{\theta}_{\hat{\bm{\epsilon}}}$, expressed as:
$\begin{split}\mathcal{L}_{kl}&\coloneqq D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}_{m}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}))\\ &=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{{\bm{z}}}}}A(\bm{\theta}_{\hat{{\bm{z}}}})-\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\nabla_{\bm{\theta}_{\hat{\bm{\epsilon}}}}A(\bm{\theta}_{\hat{\bm{\epsilon}}}).\end{split}$ (18)
Because $D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))$ is followed as:
@@
$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}&=\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\cdot\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}[\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\mathbf{T}({\bm{x}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+B({\bm{x}})]\text{d}{\bm{x}}\\ &=-A(\bm{\theta}_{\hat{{\bm{z}}}})\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})[\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\mathbf{T}({\bm{x}})+B({\bm{x}})]\text{d}{\bm{x}}\\ &=-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}B({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}},\end{split}$ (20)
and
$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}&=-A(\bm{\theta}_{\hat{\bm{\epsilon}}})\\ &+\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}B({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}.\end{split}$ (21)
$\begin{split}\therefore D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})\\ &+\bm{\theta}_{\hat{{\bm{z}}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &-\bm{\theta}_{\hat{\bm{\epsilon}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}.\end{split}$ (22)
The mean of the sufficient statistic is followed as:
$\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta})\text{d}{\bm{x}}=\frac{\partial A^{\ast}(\bm{\theta})}{\partial\bm{\theta}}\approx\frac{\partial A(\bm{\theta})}{\partial\bm{\theta}}~{}~{}~{}~{}~{}\because A^{\ast}(\bm{\theta})=\bm{\theta}^{\intercal}A^{\ast},$ (23)
where $A^{\ast}(\cdot)$ is a true log-partition function of the exponential family (ideal case of $A(\cdot)$). However, estimating $A^{\ast}$ is difficult, and there is no direct method without random samplings, such as mini-batch weighted sampling or mini-batch stratified sampling Chen et al. (2018). Then, we approximate $A^{\ast}$ to $A$, and train $A$ to be close to $A^{\ast}$. Consequently, we obtain KL divergence of the exponential family as:
@@
$\begin{split}\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}(\mathbf{x}|\bm{\theta}_{\hat{\bm{\epsilon}}})&=-Z(\bm{\theta}_{\hat{\bm{\epsilon}}})+\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{\bm{\epsilon}}})}{\partial\bm{\theta}_{\hat{\bm{\epsilon}}}}\\ &+\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})B({\bm{x}})\text{d}{\bm{x}}.\end{split}$ (25)
Therefore, the final Kullback-Leibler divergence of exponential family is followed as:
$\begin{split}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{{\bm{z}}}})}{\partial\bm{\theta}_{\hat{{\bm{z}}}}}-\bm{\theta}_{\hat{\bm{\epsilon}}}^{\intercal}\frac{\partial A(\bm{\theta}_{\hat{\bm{\epsilon}}})}{\partial\bm{\theta}_{\hat{\bm{\epsilon}}}}.\end{split}$ (26)
C.1 Integration for Multiple IPE-Transformation and EF-Conversion

### v3

@@ Appendix B Proof
By the Lemma B.1, if $q_{\phi}$ is equivariant over defined on group of symmetries $G_{I}^{J}$ and $G_{L}^{J}$, then $\psi_{G_{S}}(q_{\phi}(\cdot))$ is equivariant to symmetries in $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$ by the equivariance of $q_{\phi}$.
@@
Proof. The function $\psi_{G_{S}}(\cdot)$ is an equivariant function over group elements in $G_{S}\cap G_{L}$ by Proposition 4.1. Then, the composite function, $\psi_{G_{S}}(\cdot)$ and $q_{\phi}$, is an equivariant function of $G_{I}$ corresponding to $G_{S}\cap G_{L}$ and $G_{T}$ corresponding to $G_{S}\cap G_{L}$. Let $g_{L}^{J}$ be a group element in $G_{S}\cap G_{L}$, and $g_{I}^{J}$ is a group element in $G_{I}$ corresponding to $G_{S}\cap G_{L}$, and $g_{T}^{J}$ is a group element where corresponding to $G_{S}\cap G_{L}$ on the latent vector space transformed from the original latent vector space. Then, group element $g_{T}^{J}$ is equal to $g_{L}^{J}$:
@@
$\begin{split}&\text{then}~g_{L}^{J}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{J}\psi_{G_{S}}({\bm{z}}_{1})~(\because\hat{{\bm{z}}}_{2}=g_{T}^{J}\hat{{\bm{z}}}_{1})\\ &\Rightarrow(g_{L}^{J}-g_{T}^{J})\psi_{G_{S}}({\bm{z}}_{1})=\mathbf{0},\end{split}$ (19)
[…]
@@
Then, $\psi_{G_{S}}(g_{L}^{J}{\bm{z}}_{1})=g_{L}^{J}\psi_{G_{S}}({\bm{z}}_{1})=g_{T}^{J}\psi_{G_{S}}({\bm{z}}_{1})$. The encoder is an equivariant function over input space $\mathcal{X}$ as $q_{\phi}(g_{I}^{J}{\bm{x}}_{1})=g_{L}^{J}q_{\phi}({\bm{x}}_{1})$. Mixing two equivarience property, we can derive another equivariance relation $g_{T}^{J}\psi_{G_{S}}(q_{\phi}({\bm{x}}_{1}))=\psi_{G_{S}}(q_{\phi}(g_{I}^{J}{\bm{x}}_{1}))$ This result implies that the equivariance between input space and a latent space is preserved for $G_{S}\cap G_{L}$ if the latent vector $\bm{z}$} is transformed by $\psi_{G_{S}}.\blacksquare$
@@
We show that $\psi_{G_{S}}$ preserves equivariance between $G_{L}^{J}$ and $G_{I}^{J}$. If there exists equivariant function between input and latent vector space, there should be a group $G_{L}$ for a latent space and its corresponding group $G_{I}$ in an input space by definition of equivariance ($q_{\phi}(g_{I}x)=g_{L}q_{\phi}(x)$). In other words, $\psi_{G_{S}}(\cdot)$ guarantees to preserve the equivariance of I2L-transformation to certain symmetries in $G_{S}\cap G_{L}$ after IPE-transformation as shown in Fig. 2.
Let $P(G):=\frac{\mu(G\cap K)}{\mu(K)}$ by the assumption 4 in section 4, and $Pr(\psi_{G^{\prime}}(\cdot)\in G)=P(\psi(\cdot)\in G|\psi(\cdot)\in G^{\prime})=\frac{P(G\cap G^{\prime})}{P(G^{\prime})}$ after VAE training. Then,
Proposition B.3.
Under Assumption 4, the following probability inequality holds: $Pr(\psi_{E_{S}}(\cdot)\in G_{S})>Pr(\psi_{E_{M}}(\cdot)\in G_{S})>Pr(\psi_{M}(\cdot)\in G_{S})$, where probabilities are defined with respect to the normalized uniform measure on a compact subset $K\subset M_{n}(\mathbb{R})$.
Proof. Since $G_{S}\subset E_{S}\subset E_{M}\subset M_{n}(\mathbb{R})$,
$\Pr(\psi_{E_{S}}(\cdot)\in G_{S})=\frac{P(G_{S}\cap E_{S})}{P(E_{S})}=\frac{P(G_{S})}{P(E_{S})}=\frac{\mu(G_{S}\cap K)}{\mu(E_{S}\cap K)},\,\,\big(\because G_{S}\subset E_{S}\big)$
$\Pr(\psi_{E_{M}}(\cdot)\in G_{S})=\frac{P(G_{S}\cap E_{M})}{P(E_{M})}=\frac{P(G_{S})}{P(E_{M})}=\frac{\mu(G_{S}\cap K)}{\mu(E_{M}\cap K)},\,\,\big(\because G_{S}\subset E_{M}\big)$
$\Pr(\psi_{M}(\cdot)\in G_{S})=\frac{P(G_{S}\cap M_{n}(\mathbb{{R}}))}{P(M_{n}(\mathbb{{R}}))}=\frac{P(G_{S})}{P(M_{n}(\mathbb{R}))}=\frac{\mu(G_{S}\cap K)}{\mu(M_{n}(\mathbb{R})\cap K)},\,\,\big(\because G_{S}\subset M_{n}(\mathbb{R})\big)$
where $\mu$ is the normalized uniform measure on $K$. Because $G_{S}\subseteq E_{S}\subseteq E_{M}\subseteq M_{n}(\mathbb{R})$ and all denominators are finite by compactness of $K$, the ratios satisfy
$\frac{\mu(G_{S}\cap K)}{\mu(E_{S}\cap K)}>\frac{\mu(G_{S}\cap K)}{\mu(E_{M}\cap K)}>\frac{\mu(G_{S}\cap K)}{\mu(M_{n}(\mathbb{R})\cap K)}.$
Thus the stated inequality follows. $\blacksquare$
B.1 KL Divergence for Evidence of Lower Bound
@@
The KL divergence of Gaussian distribution (Kingma and Welling, 2013) is computed using mean and variance, which are the parameters of a Gaussian distribution. To introduce a loss as the KL divergence of Gaussian distribution, we compute KL divergence of the exponential family in Eq. 6 using the learnable parameter $T(\cdot)$ and $A(\cdot)$ with given natural parameter $\bm{\theta}_{\hat{{\bm{z}}}}$ and $\bm{\theta}_{\hat{\bm{\epsilon}}}$, expressed as:
[…]
Because $)D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}$)) is followed as:
@@
$\begin{aligned} \int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}&=\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\cdot\\ &~~~~~~~~~~~~~~[\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\mathbf{T}({\bm{x}})-A(\bm{\theta}_{\hat{{\bm{z}}}})+B({\bm{x}})]\text{d}{\bm{x}}\end{aligned}$
Expanding the Eq. 22, and $\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}=1$, then
$\begin{aligned} \int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\mathsf{d}{\bm{x}}&=-A(\bm{\theta}_{\hat{{\bm{z}}}})+\bm{\theta}_{\hat{{\bm{z}}}}^{\intercal}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &+\int^{\infty}_{-\infty}B({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}.\end{aligned}$
Expanding $\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}$ using Eq. 22 and 23 and plug into Eq. 21 with Eq. 23, we have
$\begin{split}D_{\mathrm{KL}}(f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})||f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}}))&=A(\bm{\theta}_{\hat{\bm{\epsilon}}})-A(\bm{\theta}_{\hat{{\bm{z}}}})\\ &+\bm{\theta}_{\hat{{\bm{z}}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\text{d}{\bm{x}}\\ &-\bm{\theta}_{\hat{\bm{\epsilon}}}\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\text{d}{\bm{x}}.\end{split}$ (24)
The gradient of the log-partition function with respect to the natural parameters yields the expected value of the sufficient statistics: $\nabla_{\bm{\theta}}A(\bm{\theta})=\mathbb{E}_{p(x;\bm{\theta})}[T(x)]$ (Bishop, 2006). Then the mean of the sufficient statistic is followed as:
$\int^{\infty}_{-\infty}T({\bm{x}})f_{{\bm{x}}}({\bm{x}}|\bm{\theta})\text{d}{\bm{x}}=\frac{\partial A^{\ast}(\bm{\theta})}{\partial\bm{\theta}}\approx\frac{\partial A(\bm{\theta})}{\partial\bm{\theta}}~~~~~\because A^{\ast}(\bm{\theta})=\bm{\theta}^{\intercal}A^{\ast},$
where $A^{\ast}(\cdot)$ is a true log-partition function of the exponential family (ideal case of $A(\cdot)$). However, estimating $A^{\ast}$ is difficult, and there is no direct method without random samplings, such as mini-batch weighted sampling or mini-batch stratified sampling (Chen et al., 2018). Then, we approximate $A^{\ast}$ to $A$, and train $A$ to be close to $A^{\ast}$. Plugging Eq. 25 into Eq. 23 then we have
@@
Applying same work to $\int^{\infty}_{-\infty}f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{{\bm{z}}}})\log f_{{\bm{x}}}({\bm{x}}|\bm{\theta}_{\hat{\bm{\epsilon}}})\mathsf{d}{\bm{x}}$ and plug into Eq. 21 with Eq. 26, then the final Kullback-Leibler divergence of exponential family is followed as:
[…]
Appendix C Objective Function